# Batched Line Search Strategy for Navigating through Barren Plateaus in Quantum Circuit Training

### v3

@@ 5 Results
$CNOT$ $CNOT$ $CNOT$ $\overline{T}\,[s]$ $CNOT$ $\overline{T}\,[s]$
4gt5_76 $5$ $46$ $529$ $24$ $1711$ $22$ $79$
4gt10-v1_81 $5$ $66$ $372$ $39$ $65737$ $37$ $2245$
one-two-three-v1_99 $5$ $59$ $302$ $45$ $80390$ $35$ $4693$
one-two-three-v0_98 $5$ $65$ $213$ $61$ $175994$ $42$ $20313$
one-two-three-v2_100 $5$ $32$ $502$ $37$ $5141$ $27$ $2687$
4mod7-v1_96 $5$ $72$ $150$ $33$ $10255$ $26$ $2951$
aj_e11_165 $5$ $69$ $337$ $36$ $15585$ $23$ $2823$
alu-v2_32 $5$ $72$ $469$ $41$ $33820$ $31$ $1639$
alu-v4_36 $5$ $51$ $193$ $40$ $11090$ $28$ $2106$
4gt5_77 $5$ $58$ $338$ $19$ $2855$ $17$ $483$
4gt12-v0_86 $6$ $116$ $1837$ $40$ $8174$
4gt12-v0_87 $6$ $112$ $1837$ $47$ $36$ $8598$
4gt12-v0_88 $6$ $86$ $1837$ $44$ $36$ $6577$
@@
Table 1: $CNOT$ gate count comparison of deep $4-6$-qubit circuits obtained by decomposing and optimizing unitaries taken from the online database [80]. The individual columns label the name of the circuit, the number of qubits in the circuit, the initial $CNOT$ gate count of the circuit, the $CNOT$ gate count of the quantum circuit synthesized by QISKIT transpile function. The last four columns show the $CNOT$ count and average execution times obtained from Refs. [57] and [79], and the same kind of results obtained with the new evolutionary algorithm with the SQUANDER package. The final error $f$ of the approximation calculated via Eq. (18) was less than $10^{-8}$ in each decomposition corresponding. We performed the benchmark calculations on a computing server equipped with $32$-Core AMD EPYC 7542 Processor (providing $64$ threads with multi-threading) and with $128$GB of memory.
@@
Our numerical experiences revealed that quantum gate synthesis provides a perfect opportunity to explore problems where the presence of local minima in the optimization landscape causes traditional optimizers to fail, even at smaller scales. In our experiments, we observed that employing the developed evolutionary algorithm with multiple agents is indeed necessary to synthesize deep circuits while achieving a significant reduction in gate count. We executed $64$ concurrent agents and synchronizing their evolution after $200-2000$ optimization steps. To reduce the computational footprint, we limited the randomly chosen set of parameters to contain only a single parameter. Nevertheless, the developed evolutionary algorithm demonstrated exceptional efficiency in solving the addressed problems. For optimal performance, we combined the evolutionary algorithm with the BFGS strategy, switching to the latter when the cost function was reduced down to $10^{-3}$ by the evolutionary algorithm. The results of our numerical experiments are presented in Table. 1. Comparing these results with previous findings, we can draw two conclusions. Firstly, the average execution time was reduced by an order of magnitude in all addressed use cases where comparisons could be made. Secondly, the evolutionary algorithm exhibited remarkable stability and efficiency in solving the optimization problem with a precision of $10^{-8}$.
@@ Discussion
We applied the optimization strategy to VQE simulations on randomly generated Heisenberg XXX and Sachdev–Ye–Kitaev (SYC) models, all of which were previously identified to exhibit barren plateaus (BPs) during the VQE process. We increased the depth of the training circuits to $p=1000$ entangling layers with $16$ qubits. The developed optimization strategy, utilizing the gradient-free search direction method, consistently decreased the VQE energy over thousands of iterations without getting trapped by BPs. We demonstrated these outstanding solver capabilities by training quantum circuits containing up to $15000$ entangling gates.
@@
Combining the described probabilistic optimization strategy with machine learning practices has the potential to further improve its performance. Replacing the probability-based selection with machine learning inference could significantly reduce the number of redundant evaluations of the cost function during circuit training. We leave the exploration of this possibility for future research projects.
@@ Acknowledgments
Founding: This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004), by the ÚNKP-23-5 New National Excellence Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund, and by the Hungarian Scientific Research Fund (OTKA) Grants No. K134437 and FK135220. We aslo acknowledge the support from the QuantERA II project HQCC-101017733. RP. acknowledge support from the Hungarian Academy of Sciences through the Bolyai János Stipendium (BO/00571/22/11) as well. We acknowledge the computational resources provided by the Wigner Scientific Computational Laboratory (WSCLAB) (the former Wigner GPU Laboratory).
Data and materials availability: The compiled quantum circuits reported in this work and the VQE trained circuits are publicly available within the SQUANDER package in QASM and other binary formats at GitHub repository[48]

### v4

@@ 5 Results
CNOTCNOT CNOTCNOT CNOTCNOT $\overline{T}\,[s]$ CNOTCNOT $\overline{T}\,[s]$
4gt5_76 55 $46$ $529$ $24$ $1711$ $22$ $79$
4gt10-v1_81 55 $66$ $372$ $39$ $65737$ $37$ $2245$
one-two-three-v1_99 55 $59$ $302$ $45$ $80390$ $35$ $4693$
one-two-three-v0_98 55 $65$ $213$ $61$ $175994$ $42$ $20313$
one-two-three-v2_100 55 $32$ $502$ $37$ $5141$ $27$ $2687$
4mod7-v1_96 55 $72$ $150$ $33$ $10255$ $26$ $2951$
aj_e11_165 55 $69$ $337$ $36$ $15585$ $23$ $2823$
alu-v2_32 55 $72$ $469$ $41$ $33820$ $31$ $1639$
alu-v4_36 55 $51$ $193$ $40$ $11090$ $28$ $2106$
4gt5_77 55 $58$ $338$ 1919 $2855$ $17$ $483$
4gt12-v0_86 66 $116$ $1837$ - - $40$ $8174$
4gt12-v0_87 66 $112$ $1837$ $47$ - $36$ $8598$
4gt12-v0_88 66 $86$ $1837$ $44$ - $36$ $6577$
@@
Table 1: CNOTCNOT gate count comparison of deep 464-6-qubit circuits obtained by decomposing and optimizing unitaries taken from the online database [82]. The individual columns label the name of the circuit, the number of qubits in the circuit, the initial CNOTCNOT gate count of the circuit, the $CNOT$ gate count of the quantum circuit synthesized by QISKIT transpile function. The last four columns show the $CNOT$ count and average execution times obtained from Refs. [57] and [81], and the same kind of results obtained with the new evolutionary algorithm with the SQUANDER package. The final error $f$ of the approximation calculated via Eq. (18) was less than $10^{-8}$ in each decomposition corresponding. We performed the benchmark calculations on a computing server equipped with $32$-Core AMD EPYC 7542 Processor (providing $64$ threads with multi-threading) and with 128128GB of memory.
@@
Our numerical experiences revealed that quantum gate synthesis provides a perfect opportunity to explore problems where the presence of local minima in the optimization landscape causes traditional optimizers to fail, even at smaller scales. In our experiments, we observed that employing the developed evolutionary algorithm with multiple agents is indeed necessary to synthesize deep circuits while achieving a significant reduction in gate count. We executed $64$ concurrent agents and synchronizing their evolution after 2002000200-2000 optimization steps. To reduce the computational footprint, we limited the randomly chosen set of parameters to contain only a single parameter. Nevertheless, the developed evolutionary algorithm demonstrated exceptional efficiency in solving the addressed problems. For optimal performance, we combined the evolutionary algorithm with the BFGS strategy, switching to the latter when the cost function was reduced down to $10^{-3}$ by the evolutionary algorithm. The results of our numerical experiments are presented in Table. 1. Comparing these results with previous findings, we can draw two conclusions. Firstly, the average execution time was reduced by an order of magnitude in all addressed use cases where comparisons could be made. Secondly, the evolutionary algorithm exhibited remarkable stability and efficiency in solving the optimization problem with a precision of $10^{-8}$.
@@ Discussion
We applied the optimization strategy to VQE simulations on randomly generated Heisenberg XXX and Sachdev–Ye–Kitaev (SYC) models, all of which were previously identified to exhibit barren plateaus (BPs) during the VQE process. We increased the depth of the training circuits to $p=1000$ entangling layers with $16$ qubits and to $p=500$ entangling layers with $21$ qubits . The developed optimization strategy, utilizing the gradient-free search direction method, consistently decreased the VQE energy over thousands of iterations without getting trapped by BPs. We demonstrated these outstanding solver capabilities by training quantum circuits containing up to $15000$ entangling gates.
@@
Combining the described probabilistic optimization strategy with machine learning practices has further potential to further improve its performance. Replacing the probability-based selection with machine learning inference could significantly reduce the number of redundant evaluations of the cost function during circuit training. We leave the exploration of this possibility for future research projects.
This work focuses on the developing and testing optimization strategies in an ideal, noise-free environment to understand their theoretical behavior before tackling the effects of hardware imperfections during the circuit training. We plan an extensive study on optimal circuit shot counts, which we expect to vary throughout the training and depend on other hyperparameters such as the batch size. Additional effects like qubit relaxation or amplitude/phase damping may also impact performance, and we see their investigation as a substantial project in its own right. Such future work will build on the present theoretical foundation to prepare the method for noisy, real-world quantum hardware.
@@ Acknowledgments
Founding: This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004), by the ÚNKP-23-5 New National Excellence Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund, and by the Hungarian Scientific Research Fund (OTKA) Grants No. K134437 and FK135220. We also acknowledge the support from the QuantERA II project HQCC-101017733 and from the Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA via the project 101113946 OpenSuperQPlus100. RP. acknowledge support from the Hungarian Academy of Sciences through the Bolyai János Stipendium (BO/00571/22/11) as well. We acknowledge the computational resources provided by the Wigner Scientific Computational Laboratory (WSCLAB) (the former Wigner GPU Laboratory).